# Homogenization Effects of Large Language Models on Human Creative Ideation

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 16th ACM Conference on Creativity & Cognition; June 24–26; Chicago, ILprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanitiesccs: Computing methodologies Natural language processing
@@ 4. RQ4
    Beyond originality, do ChatGPT and non-ChatGPT users differ in terms of other facets of creativity, such as fluency, flexibility, and elaboration? (A: ChatGPT normal-$\rightarrow$ higher fluency, flexibility, and elaboration)
@@ 2.1. Prior Studies of Homogenization
While not addressing homogenization effects directly, several other recent studies examine how LLMs change the writing of humans who use them for writing support. Lee et al. (Lee et al., 2022) find that LLM support tends to increase the diversity of a user’s vocabulary but may reduce their feelings of ownership for the text they produce. Jakesch et al. (Jakesch et al., 2023) find that an opinionated LLM-based CST influences the opinions expressed by its users in argumentative writing. Similarly, Bhat et al. (Bhat et al., 2023) find that LLM-supplied next-phrase suggestions may alter the form and content of a human user’s writing even when the user dislikes these suggestions. Potentially explaining these results, Roemmele (Roemmele, 2021) proposes and provides evidence for an “inspiration through observation” model of how human writers are influenced by LLMs: observation of LLM-generated text shapes the ideas expressed by the LLM’s user even when the user does not incorporate LLM-generated text directly into their writing.
@@ 3.1.1. Recruitment
Participants were recruited from academic mailing lists, forums, and solicitations posted by the experimenters on social media. Our study protocol and recruitment materials were approved by the Santa Clara University IRB. Participation was incentivized with a \$17.50 gift card for a one hour session. Participants were required to have a stable Internet connection, a device capable of screen-sharing, and access to quiet place for the duration of the session in order to participate in the study. Of the 36 participants originally recruited, three were excluded from all analyses for failure to follow direction (e.g. not using ChatGPT when prompted to do so).

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Creativity and Cognition; June 23–26, 2024; Chicago, IL, USAbooktitle: Creativity and Cognition (C&C ’24), June 23–26, 2024, Chicago, IL, USAdoi: 10.1145/3635636.3656204isbn: 979-8-4007-0485-7/24/06ccs: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanitiesccs: Computing methodologies Natural language processing
@@ 4. RQ4
    Beyond originality, do ChatGPT and non-ChatGPT users differ in terms of other facets of creativity, such as fluency, flexibility, and elaboration? (A: ChatGPT $\rightarrow$ higher fluency, flexibility, and elaboration)
@@ 2.1. Prior Studies of Homogenization
While not addressing homogenization effects directly, several other recent studies examine how LLMs change the writing of humans who use them for writing support (Lee et al., 2024). Lee et al. (Lee et al., 2022) find that LLM support tends to increase the diversity of a user’s vocabulary but may reduce their feelings of ownership for the text they produce. Jakesch et al. (Jakesch et al., 2023) find that an opinionated LLM-based CST influences the opinions expressed by its users in argumentative writing. Similarly, Bhat et al. (Bhat et al., 2023) find that LLM-supplied next-phrase suggestions may alter the form and content of a human user’s writing even when the user dislikes these suggestions. Potentially explaining these results, Roemmele (Roemmele, 2021) proposes and provides evidence for an “inspiration through observation” model of how human writers are influenced by LLMs: observation of LLM-generated text shapes the ideas expressed by the LLM’s user even when the user does not incorporate LLM-generated text directly into their writing.
@@ 3.1.1. Recruitment
Participants were recruited from academic mailing lists, forums, and solicitations posted by the experimenters on social media. Our study protocol and recruitment materials were approved by the Santa Clara University IRB. Participation was incentivized with a \$17.50 gift card for a one hour session. Participants were required to have a stable Internet connection, a device capable of screen-sharing, and access to a quiet place for the duration of the session in order to participate in the study. Of the 36 participants originally recruited, three were excluded from all analyses for failure to follow direction (e.g., not using ChatGPT when prompted to do so).
@@ 5.4. Mitigating Homogeneity of LLM Outputs
5.5. Implications for LLM-Based CST Design
Many of our findings follow naturally from the fact that an LLM will tend to produce similar outputs in response to similar inputs, even when those inputs come from different users. This property of LLMs is often desirable, especially when the goal is to build systems that respond reliably to queries in a particular domain. However, the fact that LLM outputs are strongly reflective of direct user inputs raises a key question for designers of LLM-based CSTs: when users are in the early stages of a creative process and their creative intent is still relatively undefined, how can an LLM-based tool avoid guiding them toward well-trodden ground?
A vaguely defined creative intent, expressed as (for instance) a short text string containing only a handful of words, will often tend—for level-of-detail reasons alone—to be similar to vaguely defined intents expressed by other users. For an LLM to directly transform a brief expression of intent into a substantially larger artifact (e.g., a passage of text), it will necessarily have to make a large number of creative decisions (e.g., individual word choices) on the user’s behalf, and these decisions will be made by reference to the relative likelihood of different alternatives. Brief statements of creative intent therefore yield “samey” artifacts, which exhibit a phenomenon characterized elsewhere as the dearth of the author (Kreminski, 2024). In other words, to get highly original outputs from an LLM, the user must supply the LLM with an input that contains enough detail to strongly differentiate it from other users’ inputs; the LLM itself functions most directly as a means of refracting user-supplied inputs through the lens of statistical patterns learned from its training data.
In light of this limitation, we believe that it may be necessary to move toward a view of creativity support as intent elicitation (Kreminski and Chung, 2024): progressively drawing a more and more detailed, idiosyncratic specification of creative intent out of the CST user, with the goal of getting users to think more rather than less about each creative decision that they make along the way. By scaffolding the multi-turn discovery and expression of creative intent, we hope that LLM-based CSTs can eventually operate as something like a focusing lens for the imagination: a means of gradually clarifying and magnifying the spark of originality that first emerges in the user’s head, without substituting the LLM’s implicitly statistical and majoritarian process of creative decision-making for the user’s own. Whether this design vision can be effectively realized remains to be seen, but we believe it to be an especially worthwhile goal for the next generation of AI-based CSTs to pursue.
6. Limitations
Our study was conducted in a lab context and consequently may not reflect some aspects of how people use LLMs for creativity support “in the wild”. In order to keep study sessions to a manageable length, participants were only allowed a fixed and relatively short amount of time to respond to each ideation prompt; additionally, the ideation prompts in our study were supplied by us, rather than being formulated independently by CST users. Most organic divergent ideation scenarios are not subject to these constraints.
Because our study made use of a general-purpose LLM-based chatbot (ChatGPT) rather than an LLM-based CST specifically designed to support divergent ideation, it should be noted that alternative deployments of LLMs may be able to avoid homogenizing user ideas. Padmakumar and He have found evidence that feedback-tuned LLMs may induce creative homogenization where base (non-feedback-tuned) LLMs do not (Padmakumar and He, 2023), and task-specific CST design has previously shown promise for reducing conformity effects in crowdsourced divergent ideation (Yu et al., 2014). These findings suggest that interventions at both the AI and UI layers may help to mitigate homogenization effects in future LLM-based CSTs.